A water company in jeddah wants to borrow money for a new production facility. If the company offers to repay the loan 6000$ in year 1 and amounts increasing by 12,000 each year throgh year 6 how much can the company borrow at interest rate of 10% per year ?


1
Expert's answer
2021-04-30T10:38:02-0400

Borrowing can be calculated by using he following formula.

"Borrowing=\frac{Repayment_1}{[1+interest]}+\frac{Repayment_2}{[1+interest]^2}+\frac{Repayment_3}{[1+interest]^3}+\frac{Repayment_4}{[1+interest]^4}+\frac{Repayment_5}{[1+interest]^5}+\frac{repayment_6}{[1+interest]^6}"


"=\frac{6000}{[1+0.1]}+\frac{18000}{[1+0.1]^2}+\frac{30000}{[1+0.1]^3}+\frac{42000}{[1+0.1]^4}+\frac{54000}{[1+0.1]^5}+\frac{66000}{[1+0.1]^6}"


"=142, 341.62"